OpenStudy (anonymous):

Express the complex number in trigonometric form. -4

OpenStudy (anonymous):

4(cos 270degrees + i sin 270degrees) 4(cos 90degrees + i sin 90degrees) 4(cos 0degrees + i sin 0degrees) 4(cos 180degrees + i sin 180degrees)

OpenStudy (anonymous):

@AnswerMyQuestions @mathmath333 any help?

OpenStudy (anonymous):

No @mathmath333

OpenStudy (anonymous):

I'm not good at this stuff...

OpenStudy (mathmath333):

\(-4i\)

OpenStudy (mathmath333):

\(\Huge ?\)

OpenStudy (anonymous):

It's just -4

OpenStudy (mathmath333):

-4 cannot be a complex number it is a real number

OpenStudy (mathmath333):

or better u post a screen shot of question

OpenStudy (anonymous):

Okay I'll do that

OpenStudy (anonymous):

@mathmath333 here it is

OpenStudy (mathmath333):

well i think its a typo their it should be \(\large -4i\) or something else

OpenStudy (mathmath333):

in question 4 also,it their also only digit or \(\large i\) also

OpenStudy (anonymous):

Okay well I guess and it was wrong. I don't know what the question meant but I got it wrong anyway. Thanks for your help.

OpenStudy (anonymous):

A medal for your troubles

OpenStudy (mathmath333):

here is a solution if it is \(\large -4i\) \(\large \tt \begin{align} \color{black}{z=a+bi\\~\\ r=|z|=\sqrt{a^2+b^2}\\~\\ \alpha=tan^{-1}(\dfrac{b}{a})\\~\\ trig form \\~\\ z=r(cos\alpha+isin\alpha)\\~\\ z=0-4i\\~\\ r=\sqrt{(-4)^2}=4\\~\\ \alpha=tan^{-1}\dfrac{-4}{0}=tan^{-1}(-\infty)=-\dfrac{\pi}{2}\\~\\ z=4[cos(-90)+isin(-90)]\\~\\ z=4[(cos(270)+isin(270)] } \end{align}\) so option \(\huge A\) is correct

OpenStudy (anonymous):

That's exactly the one I guessed and it was wrong

OpenStudy (mathmath333):

thats strange!!!